OpenStudy (anonymous):

Find the inverse of g(x) = x^2 - 9. Determine whether the inverse is a function.

OpenStudy (anonymous):

a) not a function \[y= \pm \sqrt{x+9}\] b) function \[y= \pm \sqrt{x+9}\] c) not a function \[ y= \pm \sqrt{x}+9\] d) function \[ y= \pm \sqrt{x}+9\]

OpenStudy (anonymous):

not likely for the inverse to be a function since \(x^2-9\) is not "one to one"

OpenStudy (anonymous):

for example \(f(4)=16-9=7\) and also \(f(-4)=16-9=7\)

OpenStudy (anonymous):

to find the inverse, set \(x=y^2-9\) and solve for \(y\) in two steps

OpenStudy (anonymous):

\[y=\sqrt{x}+9\] and its a function